12-35
a useful lide of ine method. Compute depreciation anuary 1,2019, for $36,000 which will be depreciated using the sum-. ue of 6,000 and a useful life of three years, In 202. quipment has a residual val hine has an original total useful life of four years and $3,000 salage value Wha 20, DIEciation expense for 2020? Phelps Company purchases years-digits method. The eq equipment on January 1, 2019,for S uipment has a residual $6 36,000 which will be depreciated using the sum-of value of S6,000 and a useful life of three years. In 2020, Phelps o the straight-line method. What is depreciation expense for 2020 Brief Exercise 12-35 Computing Depreciation Expense for Change in Depreciation Method L05 years -digits r changes its depreciation method for equipment t for Change Briel Exerciee 12-36 Whitney Company purchased equipme nt on January 1,2020, for $90,000. This equipment has a useful Recording Depreciation years and a residual value of $5,000. The company uses the st Error Correction LO6 straight-line depreciation meh pany discovered that it had incorrectly recorded depreciation for 2020 as the correcting entry on January 1, 2021. or rief Exercise 12-37 Ordinary repairs of $10,000 in Jan cording Depreciation r Correction L06 ment was orisinal
Answer #1

Solution 35:

January 1, 2019
Equipment purchased 36,000
Residual Value 6,000
Useful Life 3 years
Depreciation expense for 2019 by sum of year digits method = Remaining useful life of the asset/Sum of the years digits * Depreciable Cost
Remaining usefule life of asset= 3
Sum of the years digits = 3+2+1 = 6
Depreciable cost (36,000-6,000= 30,000)
Depreciation expense (2019) 15,000
2020 Depreciation method changed to SLM
Depreciation expense in 2020 will be (36,000-6,000)/3 10,000
Note: Depreciation will change retrospectively and excess depreciation charged in year 2019 will be adjusted against the retained earnings.